Fermi National Laboratory
 

Operation Summary for May 4, 2011 - May 6, 2011

 

Monday May 4

The day shift began with Operations monitoring Tevatron (TEV) store 8709 (118.25E30) and the Recycler stash (261.6E10), with the Antiproton Source (Pbar) stacking (10.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, FESS personnel reported at 8:26 AM that they were shutting off the Booster West Gallery air handler for maintenance work; they had completed their work within 15 minutes. TEV personnel began a hollow beam study. At 1:45 PM, the P1 transfer line beam permit tripped off when a quadrupole power supply (I:Q713) began tripping off due to excessive 60-hertz problems.

During the evening shift, Operations terminated store 8709 at 4:50 PM and put the TEV into HEP shot setup. Operations disabled beam to T-978 at 4:53 PM. At 5:54 PM, a Linac Klystron RF station (KRF7) tripped off due to a LCW leak on its charging supply. Operators called in Linac personnel to investigate. Operations established store 8710 at 6:03 PM with an average initial luminosity (AIL) of 393.26E30. Linac personnel reported that they had repaired the LCW leak on KRF7 at 8:05 PM. Operations reported some short downtimes due to the MI-8 collimators, MRF5, BRF14, and a Lambertson (LAM42).

Thursday May 5

The midnight shift began with Operations monitoring store 8710 (159.93E30) and the stash (134.2E10), with Pbar stacking (11.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operators reported that a Booster vacuum pump (B:IP809) had tripped off. It would reset, but then immediately trip off again. Operators investigated and found two other downstream vacuum pumps off. Operators turned on all three pumps. Operations established beam to T-978 at 4:18 AM. Operations reported at 5:49 AM that MI-8 collimator beam positions were causing MiniBooNE trips; operators tuned beam positions.

During the day shift, Operations terminated store 8710 at 11:09 AM. Operators prepared Booster for access to replace power amplifiers for BRF8 and BRF18 and to repair the gap monitor for BRF19. Operators prepared Linac for access to replace the power amplifier socket for LRF4 and to replace beam detectors on the berm and the Booster chute. Operators prepared D0 and CDF for access for detector maintenance. Linac and Booster completed their work and Operations resumed sending beam to MiniBooNE at 2:48 PM. By 2:52 PM, Operations had established beam to NuMI, T-978, and Pbar had resumed stacking. D0 and CDF came out of access. Operators put the TEV through a dry squeeze at 3:05 PM.

During the evening shift, Operators put the TEV through a wet squeeze at 4:39 PM. Operations disabled beam to T-978 and put the TEV into shot setup at 5:07 PM. Operations established store 8714 at 6:15 PM with an AIL of 376.3E30. At 10:55 PM, MIRF7 tripped off and wouldn’t reset. Operators bypassed the station.

Friday May 6

The midnight shift began with Operations monitoring store 8714 (149.11E30) and the stash (156.7E10), with Pbar stacking (20.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 12:48 AM, the MI gap clearing kicker tripped off; they reset. Operations established beam to T-978 at 4:22 AM. At 5:29 AM, a TEV sector B1 air compressor tripped off; operators reported that its air dryer was stuck venting. Operators isolated the air dryer.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. However, the Run Coordinator will allow cryo system personnel to replace a wet engine flywheel at F2 after store 8714 is terminated this morning; he will also allow a D0 access. Other work will be tied in with this 90-minute downtime. After the next store goes in, which should be around 3 PM, the Run Coordinator will allow MI personnel an access to fill their Fluorinert system and to replace a power amplifier for MIRF7. Booster personnel will conduct a study and bring BRF8 online today. MTA will take beam this afternoon.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/06/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory